### v1

# MDPs with a State Sensing Cost

Abstract
In many practical sequential decision-making problems, tracking the state of the environment incurs a sensing/communication/computation cost. In these settings, the agent’s interaction with its environment includes the additional component of deciding when to sense the state, in a manner that balances the value associated with optimal (state-specific) actions and the cost of sensing. We formulate this as an expected discounted cost Markov Decision Process (MDP), wherein the agent incurs an additional cost for sensing its next state, but has the option to take actions while remaining ‘blind’ to the system state.
We pose this problem as a classical discounted cost MDP with an expanded (countably infinite) state space. While computing the optimal policy for this MDP is intractable in general, we bound the sub-optimality gap associated with optimal policies in a restricted class, where the number of consecutive non-sensing (a.k.a., blind) actions is capped. We also design a computationally efficient heuristic algorithm based on policy improvement, which in practice performs close to the optimal policy. Finally, we benchmark against the state of the art via a numerical case study.
Machine Learning, ICML
1 Introduction
Markov Decision Processes (MDPs) constitute an important framework for capturing the sequential interaction between an agent and an adaptive environment that ‘responds’ to the actions of the agent. The classical framework of an MDP is that at each time $t,$ the agent sees the state, say $X_{t},$ of the environment, and takes an action $A_{t}.$ This action, in turn, generates a feedback signal (reward/cost) that depends on the state-action pair $(X_{t},A_{t}),$ and triggers a random, action-dependent, Markovian transition in the state.
However, in many applications, ‘seeing’ the current state involves a cost. For example:
$\bullet$ In healthcare applications, there is often a monetary and/or delay cost associated with sensing the state of the patient during an ongoing intervention; e.g., white blood cell monitoring during anti-HIV drug administration, laboratory tests for ICU patients (Ernst et al., 2006; Jha et al., 2009).
$\bullet$ Applications on mobile phones must pay an energy cost to sense the location/motion/environment of the user, which must be balanced with the goal of enhancing user experience (Wang et al., 2010). Similarly, in a wireless sensor network, sensing the network state might require turning on battery-operated (and therefore energy-constrained) sensors.
$\bullet$ In remote surveillance applications, there is a communication cost to be incurred in transmitting state information (in the form of images/video) to the controller.
$\bullet$ In distributed sensing applications, there is often an additional computational cost in aggregating the sensor readings to obtain the state of the system.
$\bullet$ In robotics, where an autonomous robot is engaged in a certain task, sensing the surroundings (i.e., the state) of the robot might induce a latency (cost) in task completion.
This motivates the incorporation of a cost of state sensing, as well as a model for opportunistic state sensing, into the MDP framework. However, such an incorporation introduces technical challenges. In particular, if the agent chooses not to sense the system state (while continuing to interact with the environment), it is left with a belief distribution over the system state. Incorporating this belief into the MDP formulation induces a state space explosion, which makes the decision problem intractable.
In this paper, we formulate a cost-sensing MDP built on top of a finite, infinite-horizon discounted cost baseline MDP. The ‘augmented’ MDP, which incorporates the state sensing cost, has a countably infinite state space. At each time/epoch, the agent must, in addition to taking an action, decide whether or not to sense the next state of the MDP. If it decides to sense, it incurs an additional state sensing cost in that epoch. There is thus a non-trivial trade-off between the cost induced by suboptimal actions (under state uncertainty) and the cost of state sensing.
While the ‘augmented’ MDP of interest admits a stationary Markov policy, it is computationally intractable given the infinite state space. The goal of this paper is to design tractable, near-optimal algorithms for solving it. Our key contributions are as follows:
$\bullet$ We analyse a sequence of truncated (and therefore finite) MDPs that restrict the number of consecutive non-state-sensing (a.k.a., blind) actions the agent can take. We provide a sufficient condition for the optimal policy under such a truncated MDP to also be optimal for the original (infinite state space) MDP of interest.
$\bullet$ We derive computable bounds on the suboptimality associated with the optimal policies corresponding to the truncated MDPs.
$\bullet$ We prove that if the state sensing cost is less than a certain computable threshold, then always sensing is optimal.
$\bullet$ We propose a policy-improvement-based heuristic algorithm, which selectively searches for improving blind action sequences, that is near-optimal in practice.
$\bullet$ Finally, we conduct an extensive numerical case study, comparing the proposed planning algorithms with the state of the art.
It is important to note that the proposed formulation can be posed as a Partially Observable Markov Decision Processes (POMDP) (indeed, we do benchmark the proposed algorithms against state of the art POMDP solvers; see Section 5). However, in doing so, one loses the specific problem structure that arises in the opportunistic state sensing formulation, which we seek to exploit here for computational tractability; POMDPs are known to be intractable in general (Papadimitriou & Tsitsiklis, 1987).
The remainder of this paper is organized as follows. After surveying the literature below, we formulate our sensing-cost-incorporated MDP in Section 2. In Section 3, we propose a heuristic algorithm for this MDP. Next, in Section 4, we prove sufficient conditions for the optimality of a certain class of policies, and bound their suboptimality when they do not meet these conditions. Numerical case studies are presented in Section 5, and we conclude in Section 6. Proofs of our analytical results have been omitted from the main body of the paper due to space constraints; these can be found in the appendix.
Related Literature: In light of the preceding discussion on POMDPs, we focus here only on the (few) papers that study an explicit cost-sensing formulation identical (or similar) to ours.
Formulations equivalent to the one in the present paper have been analysed in (Hansen, 1994; Bellinger et al., 2021; Nam et al., 2021; Krale et al., 2023); the last two references refer to this formulation as an Action-Contingent-Noiseless-Observable MDP, or ACNO-MDP. (Hansen, 1994) proposes a truncation-based approximation analogous to that in Section 4.2, except they provide to approximation guarantees. (Nam et al., 2021) and (Bellinger et al., 2021) focus on reinforcement learning (RL) (as opposed to the planning problem considered here). Specifically, Nam et al. focuses on developing RL algorithms for a fixed-horizon setting using the generic POMDP solver POMCP (Silver & Veness, 2010). On the other hand, Bellinger et al. adapts Q-learning for this setting by utilizing a statistical state estimator to achieve a “higher costed return” – for every non-state-sensing action, the subsequent state is simply sampled from the belief distribution. An $\epsilon$-greedy action is then taken based on the sampled state to update the Q-table, without leveraging any structure of the belief distribution while choosing the action. (Krale et al., 2023) proposes a policy improvement heuristic referred to as ATM and devises an RL algorithm to learn this heuristic; we contrast the heuristic proposed here to the ATM heuristic in Section 3, and also in our numerical case study in Section 5. Note that none of the above-mentioned papers focuses on the planning problem in a manner that exploits the specific structure of the MDP and from the standpoint of provable optimality/suboptimality guarantees.
A related formulation is considered in (Armstrong-Crews & Veloso, 2007), which treats “sensing” as a distinct action and applies a discount factor for its cost at each step a sensing action is taken. Aside from this distinction in the problem formulation, the JIV algorithm proposed in this paper is conceptually similar to the ATM heuristic proposed in (Krale et al., 2023).
Finally, another related formulation is analysed in (Reisinger & Tam, 2024); here, if the agent decides not to sense the state in any epoch, it is constrained to play the same action as in the previous epoch.
2 Problem Formulation
In this section, we formally define our MDP formulation with a state sensing cost. We do this by first defining a ‘standard’ discounted cost MDP that serves as our baseline; we subsequently incorporate a state sensing cost, and a protocol for opportunistic state sensing on the part of the agent, into this baseline MDP.
Baseline MDP: Consider an infinite horizon discounted cost MDP $\mathcal{M}(\mathcal{S},A,\mathcal{T},\mathcal{C},\alpha).$ Here,
  • $\mathcal{S}=\{1,2,\ldots,|\mathcal{S}|\}$ denotes the (finite) state space,
  • $A=\{1,2,\ldots,|A|\}$ denotes the (finite) action space,
  • $\mathcal{T}$ denotes the transition function (i.e., $\mathcal{T}(s,a,s^{\prime})$ denotes the probability of transitioning to state $s^{\prime}$ on taking action $a$ in state $s$),
  • $\mathcal{C}$ denotes the cost function (i.e., $\mathcal{C}(s,a)$ is the cost associated with taking action $a$ in state $s$),
  • $\alpha\in(0,1)$ denotes the discount factor.
With some abuse of notation, for $a\in A,$ we use $\mathcal{T}(a)$ and $\mathcal{C}(a)$ to denote, respectively, the $|\mathcal{S}|\times|\mathcal{S}|$ transition probability matrix, and the $|\mathcal{S}|\times 1$ (column) vector of costs, associated with the action $a.$ [^1] Implicit in this notation is the assumption that any action $a\in A$ can be taken in any state $s\in\mathcal{S}.$ Denoting the state at time $t$ by $X_{t},$ and the action at time $t$ by $A_{t},$ there is a well-established theory for characterizing and computing the optimal policy that minimizes the expected discounted cost
$\mathbb{E}\left[\sum_{t=0}^{\infty}\alpha^{t}\mathcal{C}(X_{t},A_{t})\right];$
see (Puterman, 2014) and  (Ross, 1992). We use $V^{*}$ and $Q^{*}$ to denote, respectively, the optimal value function and the optimal action-value function, corresponding to $\mathcal{M}.$ As is convention, we treat $V^{*}$ to be an $|\mathcal{S}|\times 1$ column vector, and $Q^{*}$ to be a $|\mathcal{S}|\times|A|$ matrix.
MDP with state sensing cost: We now incorporate a positive state sensing cost $k$ to the above baseline MDP. Formally, the protocol for the interaction between the agent and the environment is as follows: At time $t\geq 0,$
  • Agent takes action $A_{t},$ and commits to either sensing the state at the next time step, or not; in the former case, it is said to have made a sensing action, and in the latter case, it is said to have made a blind action.
  • Agent incurs cost $\mathcal{C}(X_{t},A_{t}),$ and an additional sensing cost $k$ in case it made a sensing action.
  • The next state $X_{t+1}$ gets chosen randomly as per $\mathcal{T}(X_{t},A_{t},\cdot).$ In case the agent made a sensing step, then $X_{t+1}$ is revealed to it. If the agent made a blind step (thereby ‘saving’ on the sensing cost), then $X_{t+1}$ is not revealed to it.
We assume that the agent knows its initial state $X_{0}.$ Note that if the agent chooses a blind action at time $t,$ it must make its next action $A_{t+1}$ without precise knowledge of the state $X_{t+1}$ of the environment. The goal of the agent is to minimize its expected discounted cost (including the state sensing cost), i.e.,
$\mathbb{E}\left[\sum_{t=0}^{\infty}\alpha^{t}\left(\mathcal{C}(X_{t},A_{t})+k\mathds{1}_{\{\text{sensing action at }t\}}\right)\right].$
In the remainder of this section, we formulate the above sequential decision problem as an MDP $\mathcal{M}_{k}$ with a countably infinite state space. In the following, we refer to states in the baseline MDP (i.e., the elements of $\mathcal{S}$) as ‘root states.’ The state space of $\mathcal{M}_{k}$ is defined as
$\mathcal{S}_{\infty}:=\mathcal{S}\cup\left[\mathcal{S}\times\left(\cup_{j=1}^{\infty}A^{j}\right)\right].$
Here, the state variable corresponds to the most recently sensed (root) state, along with the string of blind actions taken thereafter. Note that each state $\tilde{s}\in\mathcal{S}_{\infty}$ is associated with a belief distribution $\mathcal{B}(\tilde{s})\in\mathbb{R}^{1\times|\mathcal{S}|}$ over the set of root states. Specifically, for $\tilde{s}=(s,a_{1},a_{2},\ldots,a_{n}),$ where $s\in\mathcal{S}$ and $a_{i}\in A$ for $1\leq i\leq n,$
$\mathcal{B}(\tilde{s})=e_{s}\mathcal{T}(a_{1})\mathcal{T}(a_{2})\cdots\mathcal{T}(a_{n}),$
where $e_{s}$ denotes the unit row vector with the $s^{\text{th}}$ entry being one. By convention, for $\tilde{s}=s\in\mathcal{S},$ (i.e., right after a sensing action), $\mathcal{B}(\tilde{s})=e_{s}.$
Next, the action space $\mathcal{A}$ for $\mathcal{M}_{k}$ is defined as
$\mathcal{A}=A\times\{sense,blind\},$
where the second component of the action captures the decision of whether or not to sense the state at the next time step. Note that $\mathcal{A}$ is finite. We also write $\mathcal{A}=\mathcal{A}_{s}\cup\mathcal{A}_{b},$ where $\mathcal{A}_{s}=A\times\{sense\}$ denotes the set of sensing actions, and $\mathcal{A}_{b}=A\times\{blind\}$ the set of blind actions.
The cost function $\mathcal{C}_{\infty}:\mathcal{S}_{\infty}\times\mathcal{A}\rightarrow\mathbb{R}$ associated with $\mathcal{M}_{k}$ is defined as follows:
$\mathcal{C}_{\infty}(\tilde{s},(a,sense))$ $=\mathcal{B}(\tilde{s})\mathcal{C}(a)+k$
$\mathcal{C}_{\infty}(\tilde{s},(a,blind))$ $=\mathcal{B}(\tilde{s})\mathcal{C}(a)$
Note that the cost has been averaged over the belief distribution over the root states.
Finally, we define the transition probability function for $\mathcal{M}_{k}$ as $\mathcal{T}_{\infty}:\mathcal{S}_{\infty}\times\mathcal{A}\times\mathcal{S}_{\infty}\rightarrow\mathbb{R}$ as follows:
$\mathcal{T}_{\infty}(\tilde{s}_{1},(a,blind),\tilde{s}_{2})$ =$=\begin{cases}1,&\text{for }\tilde{s}_{2}=(\tilde{s}_{1},a)\\ 0,&\text{otherwise}\end{cases}$
$\mathcal{T}_{\infty}(\tilde{s}_{1},(a,sense),\tilde{s}_{2})$ =$=\begin{cases}0,&\text{for }\tilde{s}_{2}\notin\mathcal{S}\\ \mathcal{B}(\tilde{s})\mathcal{T}(a)e_{\tilde{s}_{2}}^{T},&\text{for }\tilde{s}_{2}\in\mathcal{S}\end{cases}$
The MDP $\mathcal{M}_{k},$ which captures opportunistic state sensing and a sensing cost, can now be defined formally using $(\mathcal{S}_{\infty},\mathcal{A},\mathcal{T}_{\infty},\mathcal{C}_{\infty},\alpha)$. Given that the state space of this MDP is countable and its action space is finite, there exists an optimal stationary policy (see (Puterman, 2014) and  (Ross, 1992)); however, this optimal policy is not amenable to an exact computation given the infinite state space. In Section 3, we propose a heuristic based on selective policy improvement, and in Section 4, we propose iterative schemes for computing an optimal (or near-optimal) policy via a truncation of the state space.
3 Heuristic Algorithm
In this section, we introduce the Selective Policy Improvement (SPI) heuristic for our opportunistic state sensing MDP. This heuristic is stated formally as Algorithm 1.
Setting up our notation, define $V_{\mathcal{M}_{k}}^{\pi}\in\mathbb{R}^{|\mathcal{S}|\times 1}$ as the value function column vector corresponding to the policy $\pi$ for the root states. The notation $\pi^{\prime}\overset{\mathcal{S}}{\succeq}\pi$ indicates that the vector $V_{\mathcal{M}_{k}}^{\pi^{\prime}}$ is element-wise less than or equal to $V_{\mathcal{M}_{k}}^{\pi}$. Finally, $\mathsf{\bf max}\ x$ (respectively, $\mathsf{\bf min}\ x$) for a vector $x$ denotes its maximum (respectively, minimum) entry.
The SPI heuristic, initialized with a certain initial policy $\pi_{init}$ and hyperparameters $\delta$ and maxsteps, calls the $\mathsf{PolicyUpdate}$ routine (see Algorithm 2) repeatedly, until the value function improvement falls below a prescribed threshold $\delta.$ The $\mathsf{PolicyUpdate}$ routine in turn improves upon its input policy $\pi_{ref}$ as follows. For each root state, it tries to find a blind action sequence having length at most maxsteps, which improves upon $\pi_{ref}.$
While the set of all such blind action paths can be quite large, $\mathsf{PolicyUpdate}$ searches this space selectively for computational tractability. Specifically, for a root state $s,$ it considers the blind action trajectory $(a_{1},a_{2},\ldots,a_{n})$ only when for each $i,$ $(a_{1},a_{2},\ldots,a_{i})$ is an improvement over $(a_{1},a_{2},\ldots,a_{i-1})$ followed by the optimal sensing action. This allows for an efficient (and greedy) search for an improving blind action trajectory.
The preceding check is performed using the following functions: for any vector $\bar{V}\in\mathbb{R}^{|\mathcal{S}|\times 1}$, define
$V_{MS}(\mathcal{B}(\tilde{s}),\bar{V})=\min_{a\in A}\bigl{(}B(\tilde{s})\mathcal{C}(a)+\alpha B(\tilde{s})\mathcal{T}(a)\bar{V}\bigr{)}+k$
$\pi_{MS}(\mathcal{B}(\tilde{s}),\bar{V})=\arg\min_{a\in A}\bigl{(}B(\tilde{s})\mathcal{C}(a)+\alpha B(\tilde{s})\mathcal{T}(a)\bar{V}\bigr{)}$
Note that $V_{MS}(\mathcal{B}(\tilde{s}),\bar{V})$ denotes the value associated with playing the optimal sensing action under belief $\mathcal{B}(\tilde{s})$ with terminal values $\bar{V}$ (MS here stands for myopic sensing). $\pi_{MS}$ denotes the corresponding optimal sensing action. Importantly, note that the SPI heuristic only performs policy evaluations over the root states of $\mathcal{M}_{k}.$
Algorithm 1 Selective Policy Improvement (SPI)
Input: Initial policy $\pi_{init}$ , maxsteps, $\delta$
Output: $\pi^{\prime}\overset{\mathcal{S}}{\succeq}\pi_{init}$
1:  $\pi^{\prime}\leftarrow\pi_{init}$
2:  $\pi_{\text{improv}}\leftarrow\textsc{PolicyUpdate}(\pi^{\prime},\text{\emph{maxsteps}})$
3:  while $\mathsf{\bf max}\ (V_{\mathcal{M}_{k}}^{\pi^{\prime}}-V_{\mathcal{M}_{k}}^{\pi_{\text{improv}}})>\delta$ do
4:     $\pi^{\prime}\leftarrow\pi_{\text{improv}}$
5:     $\pi_{\text{improv}}\leftarrow\textsc{PolicyUpdate}(\pi^{\prime},\text{\emph{maxsteps}})$
6:  end while
Algorithm 2 $\mathsf{PolicyUpdate}$
Input: $\pi_{ref},\text{\emph{maxsteps}}$
Output: $\pi_{o}\overset{\mathcal{S}}{\succeq}\pi_{ref}$
1:  $\pi_{o}\leftarrow\pi_{ref}$
2:  for $s\in\mathcal{S}$ do
3:     $\tilde{s}\leftarrow s$
4:     $steps\leftarrow 0$
5:     $\pi^{\prime}\leftarrow\pi_{ref}$
6:     $exploredstates\leftarrow\emptyset$
7:     while $steps\leq\text{\emph{maxsteps}}$ do
8:        $exploredstates\leftarrow exploredstates\cup\{\tilde{s}\}$
9:
$V_{blind}\leftarrow\min_{a\in\mathcal{A}}\Big{(}\mathcal{B}(\tilde{s})\mathcal{C}(a)+\alpha V_{MS}(\mathcal{B}(\tilde{s})\mathcal{T}(a),V_{\mathcal{M}_{k}}^{\pi_{ref}})\Big{)}$
10:
$a_{blind}\leftarrow\arg\min_{a\in\mathcal{A}}\Big{(}\mathcal{B}(\tilde{s})\mathcal{C}(a)+\alpha V_{MS}(\mathcal{B}(\tilde{s})\mathcal{T}(a),V_{\mathcal{M}_{k}}^{\pi_{ref}})\Big{)}$
11:        if $V_{MS}(\mathcal{B}(\tilde{s}),V_{\mathcal{M}_{k}}^{\pi_{ref}})\leq V_{blind}$ then
12:           $\pi^{\prime}(\tilde{s})\leftarrow(\pi_{MS}(\mathcal{B}(\tilde{s}),V_{\mathcal{M}_{k}}^{\pi_{ref}}(\mathcal{S})),sense)$
13:           break Exit the while loop
14:        end if
15:        $\pi^{\prime}(\tilde{s})\leftarrow(a_{blind},blind)$
16:        $steps\leftarrow steps+1$
17:        $\tilde{s}\leftarrow(\tilde{s},a_{blind})$
18:     end while
19:     if $V_{\mathcal{M}_{k}}^{\pi^{\prime}}(s)<V_{\mathcal{M}_{k}}^{\pi_{ref}}(s)$ then
20:        for $state\in exploredstates$ do
21:           $\pi_{o}(state)\leftarrow\pi^{\prime}(state)$
22:        end for
23:     end if
24:  end for
Finally, we note that in our numerical case studies, we initialize the SPI heuristic with the Always Sense (AS) policy, which takes the optimal sensing action in each belief state. The value function of the AS policy is, for $\tilde{s}\in\mathcal{S}_{\infty},$
$V_{AS}(B(\tilde{s}))$ $=\min_{a\in A}\bigl{(}B(\tilde{s})\mathcal{C}(a)+\alpha B(\tilde{s})\mathcal{T}(a)V^{*}\bigr{)}+\frac{k}{1-\alpha}$
$=\mathsf{\bf min}\ B(\tilde{s})Q^{*}+\frac{k}{1-\alpha}.$
Here, with some abuse of notation, we parameterize the value function $V_{AS}$ by the belief vector of state $\tilde{s},$ rather than by $\tilde{s}$ directly. The action corresponding to the AS policy is thus $(\pi_{AS}(\tilde{s}),sense),$ where
$\pi_{AS}(\tilde{s})=\mathsf{\bf argmin}\ B(\tilde{s})Q^{*}.$
Here, $\mathsf{\bf argmin}\ x$ for a row vector $x$ denotes the column index corresponding to its minimum entry. It is important to note that AS policy $\pi_{AS}$ agrees with the optimal policy $\pi^{*}$ associated with the baseline MDP $\mathcal{M}$ over root states. In Section 4.1, we show that the AS policy is also optimal for $\mathcal{M}_{k}$ when the sensing cost $k$ is small.
It is instructive at this point to compare the SPI heuristic with the ATM heuristic proposed in (Krale et al., 2023). The latter is more restricted in its search for good blind actions; in any belief state $\tilde{s},$ it seeks to improve upon $\pi_{AS}$ by comparing the actions $(\pi_{AS}(\tilde{s}),\textit{blind})$ and $(\pi_{AS}(\tilde{s}),\textit{sense}).$
4 Solving $\mathcal{M}_{k}$ with Provable Guarantees
In this section, we describe conditions and approaches that enable provable optimality/suboptimality guarantees for the MDP $\mathcal{M}_{k}.$ First, we provide a sufficient condition for always sensing to be optimal. Subsequently, we analyse a truncated version of $\mathcal{M}_{k}$ and provide sufficient conditions for the optimal solution of this (finite) truncated MDP to agree with that of $\mathcal{M}_{k}.$
4.1 Optimality of always sensing
The following theorem shows that if the state sensing cost $k$ is smaller than a certain threshold, then it is optimal to always sense the state.
Theorem 4.1.
If
$k<\alpha\ \mathsf{\bf min}\min_{a_{1},a_{2}\in A}\left[\mathcal{T}(a_{1})\left(Q^{\ast}(a_{2})-V^{\ast}\right)\right],$
then the AS policy defined in Section 3 is optimal for the MDP $\mathcal{M}_{k}.$
The threshold on state sensing cost in Theorem 4.1 may be interpreted as follows. It is strictly positive if and only if, for any action $a_{1}$ taken in any root state $j,$ there does not exist an action $a_{2}$ that is optimal in the baseline MDP on all states that lie in the belief support.
4.2 Analysis via state space truncation
We now consider a class of finite MDPs obtained via state space truncation of $\mathcal{M}_{k}.$ Specifically, the truncation is parameterized by $n\geq 0,$ the maximum number of consecutive blind actions the agent is permitted to take.
Formally, the truncated MDP, denoted by $\mathcal{M}_{k,n}$ is defined as follows. The state space is given by
$\mathcal{S}_{n}:=\mathcal{S}\cup\left[\mathcal{S}\times\left(\cup_{j=1}^{n}A^{j}\right)\right].$
Note that for $n\geq 1,$ $\mathcal{S}=\mathcal{S}_{0}\subset\mathcal{S}_{n}\subset\mathcal{S}_{n+1}\subset\mathcal{S}_{\infty}.$ We find it convenient to categorize the states of $\mathcal{S}_{n}$ into ‘layers’ as follows: Let $\mathcal{L}_{0}$ be the set of root states (the $0^{\text{th}}$ layer), which correspond to the states where the agent knows its current state precisely. Next, we define $\mathcal{L}^{j}_{m}$ as the set of ‘descendants’ of the root state $j$ in the $m^{\text{th}}$ layer, i.e., the set of states corresponding to playing $m$ successive blind steps starting from the root state $j$. More formally,
$\mathcal{L}^{j}_{m}=\{\tilde{s}\in\mathcal{S}_{\infty}\mid\tilde{s}=(j,a_{1},\ldots,a_{m}),\text{ where }a_{1},\ldots,a_{m}\in A\}.$
Finally, $\mathcal{L}_{m}$ defines the $m^{\text{th}}$ layer, defined as the union of sets $\mathcal{L}^{j}_{m}$ over all root states $j$ in $\mathcal{L}_{0}$, i.e., $\mathcal{L}_{m}=\bigcup_{j\in\mathcal{L}_{0}}\mathcal{L}^{j}_{m}.$ Note that $\mathcal{S}_{n}=\cup_{m=0}^{n}\mathcal{L}_{m}.$ Figure 1 provides an illustration of this layered view of the state space $\mathcal{S}_{n}$ for the special case of a two-state ($\mathcal{S}=\{0,1\}$), two-action ($A=\{L,R\}$) baseline MDP.
The action space $A_{n}$ of $\mathcal{M}_{k,n}$ is simply $A_{\infty}$, whereas the transition function $\mathcal{T}_{n}:\mathcal{S}_{n}\times A_{n}\times\mathcal{S}_{n}\rightarrow\mathbb{R}$ and the cost function $\mathcal{C}_{n}:\mathcal{S}_{n}\times A_{n}\rightarrow\mathbb{R}$ are given by:
$\mathcal{T}_{n}(\tilde{s}_{1},(a,\cdot),\tilde{s}_{2})$ =$=\begin{cases}\mathcal{T}_{\infty}(\tilde{s}_{1},(a,\cdot),\tilde{s}_{2}),&\text{if }\tilde{s}_{1}\notin\mathcal{L}_{n}\\ \mathcal{T}_{\infty}(\tilde{s}_{1},(a,sense),\tilde{s}_{2}),&\text{otherwise}\end{cases}$
$\mathcal{C}_{n}(\tilde{s},(a,\cdot))$ =$=\begin{cases}\mathcal{C}_{\infty}(\tilde{s},(a,\cdot)),&\text{if}\tilde{s}\notin\mathcal{L}_{n}\\ \mathcal{C}_{\infty}(\tilde{s},(a,sense)),&\text{otherwise}\end{cases}$
Note that the transition and cost functions in $\mathcal{M}_{k,n}$ agree with those in $\mathcal{M}_{k},$ except on states at the $n^{\text{th}}$ layer, where state sensing is enforced.
Since $\mathcal{M}_{k,n}$ is a finite MDP, it admits an exact computation of its optimal policy $\pi^{*}_{\mathcal{M}_{k,n}}$ and optimal value function $V^{*}_{\mathcal{M}_{k,n}}.$ Of course, the complexity of this computation grows exponentially in $n,$ so this is only feasible for small values of $n.$
In the remainder of this section, we relate the solutions of the (finite, and therefore ‘tractable’) truncated MDPs $\{\mathcal{M}_{k,n}\}$ to one another, and to the solution of $\mathcal{M}_{k}.$
Our first result bounds the suboptimality induced by the aforementioned state space truncation.
Theorem 4.2.
$V^{*}_{\mathcal{M}_{k,N}}(j)-V^{*}_{\mathcal{M}_{k}}(j)\leq\frac{\alpha^{N}k}{1-\alpha}\quad\forall j\in\mathcal{S},N\geq 0.$
Note that using the above result, one can determine a suitable truncation depth $N$ given a suboptimality tolerance, without having to solve $\mathcal{M}_{k,N}$ first. (A more refined suboptimality bound, expressed in terms of the solution of $\mathcal{M}_{k,N}$ is provided later in Theorem 4.4.)
Our next result provides a necessary and sufficient condition for the optimal policy for $\mathcal{M}_{k,n}$ to also be optimal for $\mathcal{M}_{k,n+1}.$ For $s\in\mathcal{S}$ and $a_{1},a_{2},\ldots,a_{m}\in A$, define
$Z((s,a_{1},a_{2},\ldots,a_{m})):=\mathcal{C}(s,a_{1})\\ +\sum_{i=1}^{m-1}\alpha^{i}\mathcal{B}((s,a_{1},a_{2},\ldots,a_{i}))\mathcal{C}(s,a_{i+1})$
as the average cumulative discounted cost incurred in reaching the state $\tilde{s}=(s,a_{1},a_{2},\ldots,a_{m})$ from its root state $s$ (by taking a sequence of blind steps).
[…]
Figure 1: State space of $\mathcal{M}_{k,2}$ for a 2-state 2-action baseline MDP $\mathcal{M}$
Lemma 4.3.
Fix $N\geq 0.$
$V^{*}_{\mathcal{M}_{k,N}}(j)=V^{*}_{\mathcal{M}_{k,N+1}}(j)$ for all root states $j\in\mathcal{S}$ if and only if
$Z(i)+\alpha^{N+1}$ $\min_{a}\Bigl{(}\mathcal{B}(i)\mathcal{C}(a)+k+\alpha\mathcal{B}(i)\mathcal{T}(a)V^{*}_{\mathcal{M}_{k,N}}\Bigr{)}$
$\quad\quad\geq V^{*}_{\mathcal{M}_{k,N}}(j)\quad\forall j\in\mathcal{S},i\in\mathcal{L}^{j}_{N+1}.$
Note that $V^{*}_{\mathcal{M}_{k,N}}(j)=V^{*}_{\mathcal{M}_{k,N+1}}(j)$ for all $j\in\mathcal{S}$ implies that the optimal stationary policy $\pi^{*}_{M_{k,N}}$ for $\mathcal{M}_{k,N}$ is also optimal for $\mathcal{M}_{k,N+1}$ when starting in a root state (i.e., with knowledge of the starting state). However, it is important to note that this condition does not imply that $\pi^{*}_{M_{k,N}}$ is optimal for $\mathcal{M}_{k}$ when starting in a root state. Indeed, we provide a counter-example in Section 5 to demonstrate this. One needs a stricter condition on $\mathcal{M}_{k,N}$ to conclude that $\pi^{*}_{M_{k,N}}$ is optimal for $\mathcal{M}_{k};$ this is the focus of our next result.
Define
$V_{AS;0}(\tilde{s}):=\mathsf{\bf min}\ B(\tilde{s})Q^{*}.$
This is simply the value function corresponding to the AS policy introduced in Section 3, assuming zero sensing cost. This means $V_{AS;0}$ provides a lower bound on the optimal value function for $\mathcal{M}_{k}.$
Theorem 4.4.
Fix $N\geq 0.$ If
$Z(i)+\alpha^{N+1}V_{AS;0}(i)\geq V^{*}_{\mathcal{M}_{k,N}}(j)\quad\forall j\in\mathcal{S},\ i\in\mathcal{L}^{j}_{N+1},$ (1)
then the optimal stationary policy $\pi^{*}_{\mathcal{M}_{k,N}}$ of $\mathcal{M}_{k,N}$ is optimal for $\mathcal{M}_{k}$ when starting at any root state.
If (1) does not hold,
$V^{*}_{\mathcal{M}_{k,N}}(s)-V^{*}_{\mathcal{M}_{k}}(s)\leq\epsilon_{N}\quad\forall s\in\mathcal{S},$ (2)
where
$\epsilon_{N}:=\max_{j\in S}\left[V^{*}_{\mathcal{M}_{k,N}}(j)-\min_{i\in\mathcal{L}^{j}_{N+1}}\left(Z(i)+\alpha^{N+1}V_{AS;0}(i)\right)\right].$
Theorem 4.4 provides a sufficient condition (1) for the optimal policy for $\mathcal{M}_{k,n}$ to also be optimal for $\mathcal{M}_{k}$ (assuming the starting state is a root state). Even if this condition is violated, Theorem 4.4 provides a computable upper bound on the suboptimality of the policy $\pi^{*}_{\mathcal{M}_{k,N}}$ on $\mathcal{M}_{k}.$ Thus, Theorem 4.4 suggests a recipe for computing an optimal/near-optimal policy for $\mathcal{M}_{k}$: Iteratively solve $\mathcal{M}_{k,N}$ for increasing $N,$ until either (i) the condition (4.4) is satisfied, in which case the optimal policy just computed is also optimal for $\mathcal{M}_{k},$ or (ii) the suboptimality bound $\epsilon_{N}$ is acceptably small. [^2] In fact, the following stronger statement follows from the proof of Theorem 4.4: If, for any root state $j,$ it holds that $Z(i)+\alpha^{N+1}V_{AS;0}(i)\geq V^{*}_{\mathcal{M}_{k,N}}(j)\quad\forall\ i\in\mathcal{L}^{j}_{N+1},$ (3) then the optimal policy for $\mathcal{M}_{k}$ takes at most $N$ consecutive blind steps starting from $j.$ Thus, if (3) is satisfied at certain root states, one does not need to explore depths $N+1$ and beyond at these root states.
Finally, as the following lemma shows, the suboptimality bound $\epsilon_{N}$ is decreasing in $N.$
Lemma 4.5.
It always holds that $\epsilon_{N+1}\leq\epsilon_{N},$ where $\epsilon_{N}$ is as defined in the statement of Theorem 4.4. Furthermore, if $\mathcal{M}$ is irreducible, and there exists no action that is optimal for all states in the baseline MDP, then $\epsilon_{N+1}<\epsilon_{N}$ for all $N\geq|S|-2.$
5 Numerical Case Studies
In this section, we present numerical experiments that validate and also complement the analytical results in the preceding sections. We also benchmark the proposed approaches against the state of the art.
Counter-example related to Lemma 4.3: We begin with an example that demonstrates that starting at a root state, if a certain policy $\pi$ is optimal for $\mathcal{M}_{k,N}$ as well as $\mathcal{M}_{k,N+1},$ that does not guaranteee that the $\pi$ is also optimal for $\mathcal{M}_{k}.$ Consider the two-state two-action baseline MDP shown in Figure 2, with sensing cost $k=0.005$ and discount factor $\alpha=1/2.$
010.720.720.720.72, 0.0660.0660.0660.0660.9340.9340.9340.934, 0.5020.5020.5020.5020.280.280.280.28, $0.0660.066$, 0.5020.5020.5020.5020.690.690.690.69, 0.290.290.290.290.4810.4810.4810.481, 0.410.410.410.410.310.310.310.31, 0.290.290.290.290.5190.5190.5190.519, $0.41$
Figure 2: Two-state two-action MDP with actions $\{Red,Blue\}$, $k=0.005$ and $\alpha=0.5$
The optimal policy and value function corresponding to the MDPs $\mathcal{M}_{k,N}$ for different choices of $N$ are tabulated in Table 1. The optimal policy is shown as the sequence of actions to take starting at any root state, terminating in a sensing action; for example, ‘BRRRR’ means to take the sequence of blind actions ‘BRRR’ and then the sensing action ‘R.’ Note that for $N\leq 3,$ the optimal policy at root states for $\mathcal{M}_{k,N}$ is to take a sensing action. However, for $N\geq 4,$ is optimal to make a sequence of blind steps in root state $1$.
Interestingly, we find in this example that the criterion for Theorem 4.4 (see (1)) is satisfied for root state 00 at $N=2$. Therefore, it is clear at that point that the optimal policy for root state 00 will take a maximal of $2$ blind steps, and we can restrict our search for the optimal policy starting at 00 until the $2^{\text{nd}}$ layer. The same criterion is not satisfied for root state $1$ for $N\leq 6.$
$N$ $\pi^{*}_{\mathcal{M}_{k,N}}(0)$ $V^{*}_{\mathcal{M}_{k,N}}(0)$ $\pi^{*}_{\mathcal{M}_{k,N}}(1)$ $V^{*}_{\mathcal{M}_{k,N}}(1)$
0 R $0.367061$ B $0.6796465$
1 R $0.367061$ B $0.6796465$
2 R $0.367061$ B $0.6796465$
3 R $0.367061$ B $0.6796465$
4 R $0.36703456$ BRRRR $0.67958256$
5 R $0.367029$ BRRRRR $0.6795691$
6 R $0.3670226$ BRRRRRR $0.6795541$
Table 1: Change in optimal policy and value function with the number of blind steps for baseline MDP in Figure 2
Another 2-State 2-Action example: We now consider another 2-state 2-action baseline MDP example, as shown in Figure 3. We evaluate the above MDP on sensing costs $(a)\ k=0.01$ and $(b)\ k=0.25$.
$\bullet$ Applying Theorem 4.1 on the above MDP gives a sensing cost threshold of $0.05$. Hence for case $(a),$ always sensing is optimal and hence the optimal policy for state 00 is $(R,sense)$ and for $1$ is $(B,sense)$.
$\bullet$ For case $(b),$ we find that the optimal policy for both the root states remains unchanged after $N\geq 2$. It turns out that the conditions of Lemma 4.3 hold for $N=2$. However, the condition of Theorem 4.4 is satisfied for $N=4$ which suggests that the optimal policy should remain unchanged for $N\geq 4;$ see Figure 4.
$\bullet$ For case $(b),$ Optimal policies for $\mathcal{M}_{k,N}$ outperform our heuristic algorithm (see Section 3) for $N\geq 1;$ see Figure 4.
Note: For computing the lower bound on $V^{*}_{\mathcal{M}_{m}}$ in Figure 4, and in the remaining figures of this section, we have used the following bound: For any root state $j,$
$V^{*}_{\mathcal{M}_{k}}(j)\geq\min\Bigl{\{}\min_{i\in\mathcal{L}^{j}_{N+1}}\left(Z(i)+\alpha^{N+1}V_{AS;0}(i)\right),V^{*}_{\mathcal{M}_{k,N}}(j)\Bigr{.}\\ \Bigl{.}-\alpha\max_{s\in\mathcal{S}\setminus\{j\}}\left[V^{*}_{\mathcal{M}_{k,N}}(s)-\min_{r\in\mathcal{L}^{s}_{N+1}}\left(Z(r)+\alpha^{N+1}V_{AS;0}(r)\right)\right]^{+}\Bigr{\}}.$
Here, $[x]^{+}$ denotes the positive part of $x.$
010.70.70.70.7, 11110.30.30.30.3, 000.90.90.90.9, 000.20.20.20.2, 11110.70.70.70.7, 000.30.30.30.3, 11110.80.80.80.8, 11110.10.10.10.1, 00
Figure 3: A two-state two-action baseline MDP with actions $\{Red,Blue\}$ and and $\alpha=0.5$
Figure 4: Applying Theorem 4.4 on MDP in Fig 3; $k=0.25$
Benchmarking Results
Finally, we evaluated the performance of the proposed SPI heuristic against the ATM heuristic proposed in (Krale et al., 2023) (both self-implemented in Python), alongside several widely-used general-purpose offline POMDP planning algorithms, including SARSOP (Kurniawati et al., 2008), Fast Informed Bound (FIB) (Hauskrecht, 2000; Kochenderfer et al., 2022), using the POMDPs.jl (Egorov et al., 2017) ecosystem of Julia packages across various tasks in the Gymnasium environment (Towers et al., 2024). Algorithms such as Incremental Pruning (Cassandra et al., 1997) and PBVI (Pineau et al., 2003) exhibited prohibitive runtimes, failing to produce effective policies even for small state spaces, while other methods like Q-MDP (Littman et al., 1995) were consistently outperformed by our approach. Even popular online planning algorithms like PO-UCT (Silver & Veness, 2010) required significantly higher planning times to generate competitive policies
Experimental Setup: Experiments were conducted on a MacBook Air with an Apple M3 chip and 16GB of memory. Hyperparameters were set according to the defaults in POMDPs.jl, with adjustments clearly stated and made to ensure comparable performance or runtime. POMDPs were initialized with a uniform distribution over the starting states. For SARSOP, reward and sensing values were scaled by a factor of 1000 to match performance on the Frozen-Lake task, and policy computation time was increased from 1s to 100s for the Taxi task.
Frozen Lake ($|\mathcal{S}|=16/64\ \&\ |A|=4$) : In this task, the agent navigates across a frozen lake and receives a reward of $+1$ upon reaching the goal state, with a discount factor of $0.9$ (Towers et al., 2024). We consider the slippery case, where the agent moves in the intended direction with a probability of $\frac{1}{3}$, and in one of the two perpendicular directions with equal probability of $\frac{1}{3}$ for each. The experiments use the default 4x4 and 8x8 grid configurations and a customized, challenging 4x4 version; see Table 2. Our results are summarized in Table 3. Note: We state rewards rather that costs, averaged uniformly over all initial states; the last column shows the maximum computation time recorded across the different sensing cost choices.
Stochastic Taxi ($|\mathcal{S}|=500\ \&\ |A|=6$) : In this environment, the agent must pick up passengers and drop them off at the desired locations in a 5x5 grid world. We use the noisy version described in (Dietterich, 1999), where each of the four navigation actions moves the agent in the intended direction with a probability of 0.8, and in one of the two perpendicular directions with an equal probability of 0.1 for each. The agent receives reward +20 for delivering a passenger, -10 for illegal “pickup” and “drop-off” actions, and -1 per step unless another reward is triggered, with a discount factor of 0.95. The initial state is uniformly sampled from 300 valid states where the passenger is neither at their destination nor inside the taxi. Our results for this example are summarized in Table 4.
S F F F
F H F H
F F F H
H F F G
F H S F
F G H F
F H H F
F F F F
S F F F F F F F
F F F F F F F F
F F F H F F F F
F F F F F H F F
F F F H F F F F
F H H F F F H F
F H F F H F H F
F F F H F F F G
Table 2: Default and custom-hard 4x4 grids with default 8x8 grid.
We see that in nearly all cases, the proposed SPI heuristic outperforms the ATM heuristic as well as FIB; we were unable to run FIB on the Stochastic Taxi model with 500 (root) states. SPI and SARSOP are comparable in performance, with either approach outperforming the other in some examples. However, note that SPI is significantly better than SARSOP on the Stochastic Taxi model. Finally, we note that the performance of SARSOP is highly sensitive to the tuning of several ‘hard to interpret’ hyperparameters; in contrast, SPI has only two (easily interpretable) hyperparameters.
Scenario Value Function (Rewards) for Sensing Costs Time (s)
0.001 0.005 0.01 0.05
Frozen-Lake 4x4 (Default)
   SPI $\mathbf{62.42}$ $\mathbf{36.53}$ $20.99$ $\mathbf{23.08}$ $0.4$
   $\pi_{\text{ATM}}$ $\mathbf{62.42}$ $36.52$ $6.72$ $16.57$ $0.04$
   $V^{*}_{\mathcal{M}_{k,3}}$ $\mathbf{62.42}$ $\mathbf{36.53}$ $20.47$ $-28.75$ $11.5$
   SARSOP $\mathbf{62.42}$ $\mathbf{36.53}$ $20.79$ $\mathbf{23.08}$ $1.7$
   FIB ($1000$ iter) $\mathbf{62.42}$ $31.43$ $\mathbf{23.08}$ $\mathbf{23.08}$ $8.3$
Frozen-Lake 4x4 (Hard)
   SPI $\mathbf{8.95}$ $\mathbf{3.69}$ $\mathbf{1.47}$ $1.35$ $0.4$
   $\pi_{\text{ATM}}$ $8.41$ 00 00 00 $0.04$
   $V^{*}_{\mathcal{M}_{k,3}}$ $8.92$ $1.36$ $-5.75$ $-36.75$ $12$
   SARSOP $\mathbf{8.95}$ $3.66$ $1.34$ $\mathbf{1.44}$ $1.6$
   FIB ($1000$ iter) $4.44$ 00 00 00 $9.5$
Frozen-Lake 8x8
   SPI $\mathbf{3.53}$ $3.33$ $3.33$ $3.33$ $3$
   $\pi_{\text{ATM}}$ $3.29$ $3.29$ $3.29$ $3.29$ $0.25$
   $V^{*}_{\mathcal{M}_{k,3}}$ $2.72$ $-4.943$ $-13.64$ $-79.09$ $205$
   SARSOP $3.36$ $\mathbf{3.36}$ $\mathbf{3.36}$ $\mathbf{3.36}$ $2$
   FIB ($20$ iter) $3.29$ $3.29$ $3.29$ $3.29$ $475$
Table 3: Performance comparison across different scenarios with varying sensing costs; rewards are multiplicatively scaled by $10^{3}$.
Taxi Sensing Costs
0.1 0.5 1 5
SPI                 Value $\mathbf{0.911}$ $\mathbf{-0.306}$ $\mathbf{-1.598}$ $\mathbf{-9.778}$
                      Time (s) 25.2 19.7 17.8 117.8
${\pi_{\text{ATM}}}$               Value $0.908$ $-0.359$ $-2.761$ $-19.664$
                      Time (s) 2 1.7 1.8 38.4
SARSOP ($100$s) Value $-22$ $-1.2$ $-2.705$ $-20$
Table 4: Performance Comparison of Average Cumulative Reward for Stochastic Taxi
6 Concluding remarks
In this paper, we have analysed a class of MDPs with a state sensing cost. Here, the agent must, in a history-dependent, opportunistic manner, determine when to sense the state of the system/environment. While these MDPs are intractable under generic planning algorithms, we exploit the special structure of these sensing-cost MDPs to devise intelligent heuristics and truncation approaches with provable optimality/suboptimality guarantees.
At a high level, this work is related to the vast recent literature on Age of Information (AoI), where the goal is to allocate resources or incur costs so as to minimize the age (a.k.a., staleness) of the state information; see (Yates et al., 2021) for a survey. However, the AoI literature does not, as such, consider the control aspect where the goal of state estimation is actually to influence the state evolution favourably. Additionally, the AoI literature employs a universal (state-independent) age/staleness penalty; in practice, one would expect that the agent would be more tolerant of delayed state information in certain states than others. The present formulation seeks to formally capture this trade-off.
Appendix A Proofs
Contents
A.1 Proof of Theorem 4.1
Proof.
Let $\pi$ be the AS policy for the MDP $\mathcal{M}_{k}$, then for all states $\tilde{s}\in\mathcal{S}_{\infty}$, $\pi(\tilde{s})=(\pi_{AS}(\tilde{s}),sense)$. We now apply policy improvement on $\pi$. Note that policy $\pi$ cannot be improved by any sensing action for any of the states, i.e.,
$Q^{\pi}_{\mathcal{M}_{k}}(\tilde{s},(a,sense))\geq V^{\pi}_{\mathcal{M}_{k}}(\tilde{s})\quad\forall\tilde{s}\in\mathcal{S}_{\infty},\ a\in A$
. Furthermore, if (4) holds, then policy $\pi$ cannot be improved by taking blind actions and is therefore optimal.
$Q^{\pi}_{\mathcal{M}_{k}}(\tilde{s},(a_{1},blind))$ $\geq Q^{\pi}_{\mathcal{M}_{k}}(\tilde{s},(a_{1},sense)),$ (4)
$\quad\forall(\tilde{s},a_{1})\in\mathcal{S}_{\infty}\times A.$
Criterion (4) holds for a state-action pair $(\tilde{s},a_{1})$ if and only if
$\mathcal{B}(\tilde{s})\mathcal{C}(a_{1})+\alpha\left(k+\mathcal{B}(\tilde{s})\mathcal{T}(a_{1})\mathcal{C}(a_{2})\right)\\ +\alpha^{2}\mathcal{B}(\tilde{s})\mathcal{T}(a_{1})\mathcal{T}(a_{2})V^{\pi}_{\mathcal{M}_{k}}\\ \geq\mathcal{B}(\tilde{s})\mathcal{C}(a_{1})+k+\alpha\mathcal{B}(\tilde{s})\mathcal{T}(a_{1})V^{\pi}_{\mathcal{M}_{k}},$
where $a_{2}$ is the action taken according to $\pi$ at the state reached by taking action $a_{1}$ at state $\tilde{s}$. Moreover, for any root state $s\in\mathcal{S}$, we have
$V^{\pi}_{\mathcal{M}_{k}}(s)=V_{AS}(\mathcal{B}(s))=V^{\ast}(s)+\frac{k}{1-\alpha}.$
Substituting this and rearranging, we get
$\frac{k}{1-\alpha}+\alpha\mathcal{B}(\tilde{s})\mathcal{T}(a_{1})\mathcal{T}(a_{2})V^{\ast}\\ \geq\frac{k}{\alpha(1-\alpha)}+\mathcal{B}(\tilde{s})\mathcal{T}(a_{1})V^{\ast}-\mathcal{B}(\tilde{s})\mathcal{T}(a_{1})\mathcal{C}(a_{2}).$
Simplifying, we get
$\mathcal{B}(\tilde{s})\mathcal{T}(a_{1})\mathcal{C}(a_{2})-\mathcal{B}(\tilde{s})\mathcal{T}(a_{1})V^{\ast}\\ +\alpha\mathcal{B}(\tilde{s})\mathcal{T}(a_{1})\mathcal{T}(a_{2})V^{\ast}\geq\frac{k}{\alpha}.$
Further simplifying, we obtain
$\mathcal{B}(\tilde{s})\mathcal{T}(a_{1})(\mathcal{C}(a_{2})+\alpha\mathcal{T}(a_{2})V^{\ast}-V^{\ast})\geq\frac{k}{\alpha},$
which implies
$\frac{k}{\alpha}\leq\mathcal{B}(\tilde{s})\mathcal{T}(a_{1})(Q^{\ast}(a_{2})-V^{\ast}).$ (5)
If condition (6) holds, then (5) is satisfied for all $(\tilde{s},a_{1})\in\mathcal{S}_{\infty}\times A$, and hence the AS policy $\pi$ is optimal for $\mathcal{M}_{k}$.
$k<\alpha\ \mathsf{\bf min}\min_{a_{1},a_{2}\in A}\left[\mathcal{T}(a_{1})\left(Q^{\ast}(a_{2})-V^{\ast}\right)\right].$ (6)
A.2 Proof of Theorem 4.2
Proof.
Consider any stationary policy $\pi$ of $\mathcal{M}_{k}$ and define the set of root states $G$, such that starting from any state $j\in G$ and following $\pi$, we play at least $N+1$ consecutive blind steps. Also define $\bar{G}:=\mathcal{S}\setminus G$.
First, we show that
$V_{\mathcal{M}_{k,N}}^{*}(i)-V_{\mathcal{M}_{k}}^{\pi}(i)\leq\sum_{j\in\mathcal{L}_{0}}p_{ij}\bigl{(}V_{\mathcal{M}_{k,N}}^{*}(j)-V_{\mathcal{M}_{k}}^{\pi}(j)\bigr{)}$
$\quad\quad\forall i\in\bar{G},$ (7)
where $p_{ij}$ denotes the probability of landing in root state $j\in\mathcal{S}$ after taking the first sensing step when starting from $i$ and following $\pi$.
Next, we show that
$V_{\mathcal{M}_{k,N}}^{*}(i)-V_{\mathcal{M}_{k}}^{\pi}(i)$ $\leq\frac{\alpha^{N}k}{1-\alpha},\quad\forall i\in G$ (8)
Finally, define $G^{\prime}=\{s\mid s\in\bar{G}\text{ and }s\not\rightarrow i,\forall i\in G\}$, i.e., starting from any state $s\in G^{\prime},$ we never reach a state in $G$ under policy $\pi.$ It is easy to see that
$V_{\mathcal{M}_{k,N}}^{*}(i)-V_{\mathcal{M}_{k}}^{\pi}(i)$ $=0,\quad\forall i\in G^{\prime}.$ (9)
We now show how the the statement of the lemma follows from (7)–(9). Treat $f(i):=V_{\mathcal{M}_{k,N}}^{*}(i)-V_{\mathcal{M}_{k}}^{\pi}(i)$ as the reward in state $i$ corresponding to a Markov chain over $\mathcal{S}$ with transition probabilities $\{p_{ij}\},$ the states in $G\cup G^{\prime}$ being absorbing states. Note that (7) implies that starting in any non-absorbing state, the average reward increases with time; moreover, eventual absorption is guaranteed with probability 1. Since the reward on absorbing states is at most $\frac{\alpha^{N}k}{1-\alpha}$ (see (8) and  (9)), it follows that
$f(i)=V_{\mathcal{M}_{k,N}}^{*}(i)-V_{\mathcal{M}_{k}}^{\pi}(i)\leq\frac{\alpha^{N}k}{1-\alpha}\quad\forall\ i\in\mathcal{S}.$
This implies the statement of the lemma, taking $\pi$ to be an optimal policy under $\mathcal{M}_{k}.$ It now remains to prove (7) and (8).
To prove (7), consider the value function for any state $i\in\bar{G}$ under policy $\pi$
$V_{\mathcal{M}_{k,N}}^{\pi}(i)=$ $Z(s_{m})+\alpha^{m}\mathcal{C}_{N}(s_{m},\pi(s_{m}))$
$\quad+\alpha^{m+1}\Biggl{(}\sum_{j\in\mathcal{L}_{0}}p_{ij}V_{\mathcal{M}_{k,N}}^{\pi}(j)\Biggr{)},$
where $s_{m}\in\mathcal{L}^{i}_{m}$, $m\leq N$, is the first state from which a sensing action is taken starting from $i$ following $\pi$. Let $B_{\pi}$ denote the Bellman operator corresponding to policy $\pi$, then
$B_{\pi}^{m+1}V_{\mathcal{M}_{k,N}}^{*}(i)=$ $Z(s_{m})+\alpha^{m}\mathcal{C}_{N}(s_{m},\pi(s_{m}))$
$\quad+\alpha^{m+1}\Biggl{(}\sum_{j\in\mathcal{L}_{0}}p_{ij}V_{\mathcal{M}_{k,N}}^{*}(j)\Biggr{)}.$
Now observe that
$B_{\pi}^{m+1}V_{\mathcal{M}_{k,N}}^{*}(i)-V_{\mathcal{M}_{k}}^{\pi}(i)=$
$\quad\alpha^{m+1}\Biggl{(}\sum_{j\in\mathcal{L}_{0}}p_{ij}\bigl{(}V_{\mathcal{M}_{k,N}}^{*}(j)-V_{\mathcal{M}_{k}}^{\pi}(j)\bigr{)}\Biggr{)}$
$\implies$ $\,V_{\mathcal{M}_{k,N}}^{*}(i)-V_{\mathcal{M}_{k}}^{\pi}(i)\leq$
$\quad\alpha^{m+1}\Biggl{(}\sum_{j\in\mathcal{L}_{0}}p_{ij}\bigl{(}V_{\mathcal{M}_{k,N}}^{*}(j)-V_{\mathcal{M}_{k}}^{\pi}(j)\bigr{)}\Biggr{)}$
[…]
Note that the above inequality implies (7).
To prove (8), first note that for any state $j\in\mathcal{S}_{n}$ we have
$V_{\mathcal{M}_{k,N}}^{*}(j)\leq V_{\mathcal{M}_{k,N}}^{\pi_{AS}}(j)=V_{\mathcal{M}_{k}}^{\pi_{AS}}(j)\leq V_{\mathcal{M}_{k}}^{*}(j)+\frac{k}{1-\alpha}.$
Following similar steps as in the proof of (7), for any root state $j\in G,$
$B_{\pi}^{N}V_{\mathcal{M}_{k,N}}^{*}(j)-V_{\mathcal{M}_{k}}^{\pi}(j)\leq$
$\quad\alpha^{N}(V_{\mathcal{M}_{k,N}}^{*}(s_{N})-V_{\mathcal{M}_{k}}^{\pi}(s_{N}))\leq\frac{\alpha^{N}k}{1-\alpha},$
$\implies$ $\,V_{\mathcal{M}_{k,N}}^{*}(j)-V_{\mathcal{M}_{k}}^{\pi}(j)\leq\frac{\alpha^{N}k}{1-\alpha},$
where $s_{N}\in\mathcal{L}^{j}_{N}$, is the state reached after playing $N$ blind steps starting from $j$ following $\pi$. This establishes (8). ∎
A.3 Proof of Lemma 4.3
Define $\pi_{N+1}$ as an extension of the policy $\pi^{*}_{\mathcal{M}_{k,N}}$ for $\mathcal{M}_{k,N+1}$. Without loss of generality (W.L.O.G.), assume that $\pi^{*}_{\mathcal{M}_{k,N}}(\tilde{s})\in A_{s}$ for all states $\tilde{s}\in\mathcal{L}_{N}$. Under $\pi_{N+1}$, states $\tilde{s}\in\bigcup_{l=0,1,\ldots,N}\mathcal{L}_{l}$ (i.e., all states from layers 00 to $N$) are mapped to actions provided by the policy $\pi^{*}_{\mathcal{M}_{k,N}}$, while states $\tilde{s}\in\mathcal{L}_{N+1}$ (i.e., states in the $N+1$ layer) are assigned arbitrary actions.
Let $S^{j}_{exp}$ denote the sequence of states $s\in\mathcal{L}^{j}_{m}$ for $m\geq 0$ that are visited under $\pi^{*}_{\mathcal{M}_{k,N}}$ starting from the root state $j$ (inclusive of $j$). Define $S_{exp}\coloneqq\cup_{j}S^{j}_{exp}$. Furthermore, define
$Z_{s_{m}}(s_{T})=\sum_{i=m}^{T-1}\alpha^{i-m}\mathcal{B}(s_{i})\mathcal{C}(a_{i})$
for $10\leq m\leq T-$ 1, where the only difference from $Z(s_{T})$ is that we start from state $s_{m}$ at $t=0$ and calculate the cumulative cost to reach $s_{T}$.
$Z(i)+\alpha^{N+1}\min_{a}\Bigl{(}\mathcal{B}(i)\mathcal{C}(a)+k+\alpha\mathcal{B}(i)\mathcal{T}(a)V^{*}_{\mathcal{M}_{k,N}}\Bigr{)}\\ \quad\quad\geq V^{*}_{\mathcal{M}_{k,N}}(j)\quad\forall j\in\mathcal{S},\ i\in\mathcal{L}^{j}_{N+1}.$ (10)
We claim that (10) is a necessary and sufficient condition for the optimal actions to remain unchanged for all states $\tilde{s}\in S_{exp}$ in every step of policy iteration. This follows from the fact that there exists an improvable action at some state $\tilde{s}\in S_{exp}$ in $\mathcal{L}_{m}$ for the improved policy $\pi^{\prime}_{N+1}$ at some step of the policy iteration algorithm if and only if (11) is satisfied for some $i\in\mathcal{L}_{N+1}$ and $a\in A$.
$Z_{\tilde{s}}(i)+\alpha^{N+1-m}\left(\mathcal{B}(i)\mathcal{C}(a)+k+\alpha\mathcal{B}(i)\mathcal{T}(a)V^{*}_{\mathcal{M}_{k,N}}\right)\\ <V^{\pi^{\prime}_{N+1}}_{\mathcal{M}_{k,N}}(\tilde{s})$ (11)
By inequality (10), we have
$Z(\tilde{s})+Z_{\tilde{s}}(i)+\alpha^{N+1-m}(\mathcal{B}(i)\mathcal{C}(a)+k\\ \quad+\alpha\mathcal{B}(i)\mathcal{T}(a)V^{*}_{\mathcal{M}_{k,N}})\geq Z(\tilde{s})+V^{*}_{\mathcal{M}_{k,N}}(\tilde{s}).$
Simplifying, we get
$Z_{\tilde{s}}(i)+\alpha^{N+1-m}(\mathcal{B}(i)\mathcal{C}(a)+k\\ +\alpha\mathcal{B}(i)\mathcal{T}(a)V^{*}_{\mathcal{M}_{k,N}})\geq V^{*}_{\mathcal{M}_{k,N}}(\tilde{s}),$
which is the necessary and sufficient condition for the optimal value function of all root states to remain unchanged even when evaluated on $\mathcal{M}_{k,N+1}$.
A.4 Proof of Theorem 4.4
Proof.
Suppose that (1) holds. Fix a state $i\in\mathcal{L}^{j}_{N+1}$ and consider a policy $\pi^{ji}$ such that we traverse $i$ starting from root state $j$ by following this policy. Let $\mathcal{M}_{0}$ denote the corresponding MDP with no sensing cost. Then,
$V_{\mathcal{M}_{0}}^{\pi^{ji}}(j)$ $\geq Z(i)+\alpha^{N+1}V^{*}_{\mathcal{M}_{0}}(i),$
$V_{\mathcal{M}_{0}}^{\pi^{ji}}(j)$ $\geq Z(i)+\alpha^{N+1}\min_{a}\left(\mathcal{B}(i)\mathcal{C}(a)+\alpha\mathcal{B}(i)\mathcal{T}(a)V^{*}\right),$
$V_{\mathcal{M}_{0}}^{\pi^{ji}}(j)$ $\geq Z(i)+\alpha^{N+1}V_{AS;0}(i).$
Let $\pi^{j}_{M}$ be a policy such that, starting from root state $j$ and following $\pi^{j}_{M}$, we take $M>N$ consecutive blind steps. Note that
$V_{\mathcal{M}_{0}}^{\pi^{j}_{M}}(j)$ $\geq\min_{i\in\mathcal{L}^{j}_{N+1}}\left(Z(i)+\alpha^{N+1}V_{AS;0}(i)\right).$ (12)
W.L.O.G., assume $\pi^{*}_{\mathcal{M}_{k,N}}(\tilde{s})\in A_{s}$ for all $\tilde{s}\in\mathcal{L}_{N}$. If condition (1) holds, then
$V^{*}_{\mathcal{M}_{k,N}}(j)\leq\min_{i\in\mathcal{L}^{j}_{N+1}}\left(Z(i)+\alpha^{N+1}V_{AS;0}(i)\right)\\ \leq V_{\mathcal{M}_{0}}^{\pi_{j}^{M}}(j)\leq V_{\mathcal{M}_{k}}^{\pi_{j}^{M}}(j)\\ \implies V^{*}_{\mathcal{M}_{k}}(j)\leq V^{\pi^{*}_{\mathcal{M}_{k,N}}}_{\mathcal{M}_{k}}(j)=V^{*}_{\mathcal{M}_{k,N}}(j)\leq V_{\mathcal{M}_{k}}^{\pi_{j}^{M}}(j)$
Thus, the optimal policy for $\mathcal{M}_{k}$ takes at most $N$ consecutive blind steps starting from $j$, and consequently, when (1) holds, $V_{\mathcal{M}_{k,N}}(j)=V_{\mathcal{M}_{k}}(j)$ for all $j\in\mathcal{S}$.
Now consider a scenario where condition (1) does not hold and notice that this condition is not satisfied if and only if $\epsilon_{N}>0$. Exactly as in the proof of Theorem 4.2 (see Section A.2), for a stationary policy $\pi$ of $\mathcal{M}_{k}$, define a set of root states $G$ such that, starting from any state $j\in G$ and following $\pi$, at least $N+1$ consecutive blind steps are taken. Similarly, define $\bar{G}:=\mathcal{S}\setminus G$. We have already proved that for root states $j\in G$,
$V_{\mathcal{M}_{k}}^{\pi}(j)\geq\min_{i\in\mathcal{L}^{j}_{N+1}}\left(Z(i)+\alpha^{N+1}V_{AS;0}(i)\right)$
$\implies$ $V_{\mathcal{M}_{k,N}}^{*}(j)-V_{\mathcal{M}_{k}}^{\pi}(j)\leq\epsilon_{N},\quad\forall i\in G.$ (13)
Now, the value function for any state $l\in\bar{G}$ under policy $\pi$ can be represented as
$V_{\mathcal{M}_{k,N}}^{\pi}(l)=$ $Z(s_{m})+\alpha^{m}\mathcal{C}_{N}(s_{m},\pi(s_{m}))$
$\quad+\alpha^{m+1}\Biggl{(}\sum_{j\in\mathcal{L}_{0}}p_{lj}V_{\mathcal{M}_{k,N}}^{\pi}(j)\Biggr{)},$
where $s_{m}\in\mathcal{L}^{l}_{m}$, $m\leq N$, is the first state from which a sensing action is taken starting from $l$ following $\pi$. Let $B_{\pi}$ is the Bellman operator corresponding to policy $\pi$, then
$B_{\pi}^{m+1}V_{\mathcal{M}_{k,N}}^{*}(l)=$ $Z(s_{m})+\alpha^{m}\mathcal{C}_{N}(s_{m},\pi(s_{m}))$
$\quad+\alpha^{m+1}\Biggl{(}\sum_{j\in\mathcal{L}_{0}}p_{lj}V_{\mathcal{M}_{k,N}}^{*}(j)\Biggr{)}.$
Now observe that
$B_{\pi}^{m+1}V_{\mathcal{M}_{k,N}}^{*}(l)-V_{\mathcal{M}_{k}}^{\pi}(l)=$
$\quad\alpha^{m+1}\Biggl{(}\sum_{j\in\mathcal{L}_{0}}p_{lj}\bigl{(}V_{\mathcal{M}_{k,N}}^{*}(j)-V_{\mathcal{M}_{k}}^{\pi}(j)\bigr{)}\Biggr{)}$
$\implies$ $\,V_{\mathcal{M}_{k,N}}^{*}(l)-V_{\mathcal{M}_{k}}^{\pi}(l)\leq$
$\quad\alpha^{m+1}\Biggl{(}\sum_{j\in\mathcal{L}_{0}}p_{lj}\bigl{(}V_{\mathcal{M}_{k,N}}^{*}(j)-V_{\mathcal{M}_{k}}^{\pi}(j)\bigr{)}\Biggr{)}$
(since $V_{\mathcal{M}_{k,N}}^{*}\leq B_{\pi}^{m+1}V_{\mathcal{M}_{k,N}}^{*}$ elementwise)
where $p_{lj}$ denotes the probability of landing in root state $j\in\mathcal{S}$ after taking the first sensing step when starting from $l$ and following $\pi$. Consider $G^{\prime}=\{s\mid s\in\bar{G}\text{ and }s\not\rightarrow i,\forall i\in G\}$, i.e., starting from any state $s\in G^{\prime}$ we never reach any state $i\in G$, under policy $\pi$ on $\mathcal{M}_{k,N}$. Therefore, we have
$V_{\mathcal{M}_{k,N}}^{*}(i)-V_{\mathcal{M}_{k}}^{\pi}(i)$ $\leq\epsilon_{N},\quad\forall i\in G,$
$V_{\mathcal{M}_{k,N}}^{*}(i)-V_{\mathcal{M}_{k}}^{\pi}(i)$ $=0,\quad\forall i\in G^{\prime},$
$V_{\mathcal{M}_{k,N}}^{*}(i)-V_{\mathcal{M}_{k}}^{\pi}(i)$ $\leq\alpha^{a_{i}}\Bigl{(}\sum_{j\in\mathcal{L}_{0}}p_{ij}\bigl{(}V_{\mathcal{M}_{k,N}}^{*}(j)-V_{\mathcal{M}_{k}}^{\pi}(j)\bigr{)}\Bigr{)},$
$\quad\forall i\in\bar{G},$
where $a_{i}$’s are policy $\pi$ and root state-dependent constants, with $a_{i}\geq 1$. Identical to the argument made in Section A.2, treat $f(i):=V_{\mathcal{M}_{k,N}}^{*}(i)-V_{\mathcal{M}_{k}}^{\pi}(i)$ as the reward in state $i$ corresponding to a Markov chain over $\mathcal{S}$ with transition probabilities $\{p_{ij}\}$, where the states in $G\cup G^{\prime}$ are absorbing. Since the reward on absorbing states is at most $\epsilon_{N}$ (see (13)), it follows that $f(i)=V_{\mathcal{M}_{k,N}}^{*}(i)-V_{\mathcal{M}_{k}}^{\pi}(i)\leq\alpha\epsilon_{N}$ for all $i\in\bar{G}\setminus G^{\prime}.$ (for all non-absorbing states). This establishes  (2)
NOTE:
1.
Even if condition (3) is satisfied and the optimal policy for a root state $j$ of $\mathcal{M}_{k}$ is restricted to having a maximum of $N$ consecutive blind steps, it does not generally imply that $V_{\mathcal{M}_{k}}^{*}(j)=V_{\mathcal{M}_{k,N}}^{*}(j)$.
2. It follows from the same proof that the stronger claim below holds for any root state $j$:
$V^{*}_{\mathcal{M}_{k}}(j)\geq\min\Bigl{\{}\min_{i\in\mathcal{L}^{j}_{N+1}}\left(Z(i)+\alpha^{N+1}V_{AS;0}(i)\right),V^{*}_{\mathcal{M}_{k,N}}(j)\Bigr{.}\\ \Bigl{.}-\alpha\max_{s\in\mathcal{S}\setminus\{j\}}\left[V^{*}_{\mathcal{M}_{k,N}}(s)-\min_{r\in\mathcal{L}^{s}_{N+1}}\left(Z(r)+\alpha^{N+1}V_{AS;0}(r)\right)\right]^{+}\Bigr{\}}.$
Here, $[x]^{+}$ denotes the positive part of $x.$
Idea: Define a separate terminal value for each of the states $j\in G$, given by
$V_{\mathcal{M}_{k,N}}^{*}(j)-V_{\mathcal{M}_{k}}^{\pi}(j)\leq V_{\mathcal{M}_{k,N}}^{*}(j)\\ -\min_{i\in\mathcal{L}^{j}_{N+1}}\left(Z(i)+\alpha^{N+1}V_{AS;0}(i)\right).$
A.5 Proof of Lemma 4.5
Proof.
Let the baseline MDP $\mathcal{M}$ be defined according to the conditions specified in the lemma. For any state $\tilde{s}\in\mathcal{S}_{\infty}$, we have
$V_{AS;0}(\tilde{s})$ $=\mathcal{B}(\tilde{s})\mathcal{C}(\pi_{AS}(\tilde{s}))+\alpha\mathcal{B}(\tilde{s})\mathcal{T}(\pi_{AS}(\tilde{s}))V^{*},$ (14)
where $\mathcal{B}(\tilde{s})=\mathcal{B}(\tilde{s})\mathcal{T}(\pi_{AS}(\tilde{s}))$. We claim that for the above non-trivial MDP, for each $a\in A$, there exists a root state $r$ such that
$V^{*}(r)$ $<\mathcal{C}(r,a)+\alpha e_{r}\mathcal{T}(a)V^{*}.$ (15)
Also note that $\exists\ N^{*}$ s.t. $\forall N\geq|\mathcal{S}|-1$, every element of $\mathcal{B}(\tilde{s})$ is non-zero $\forall\tilde{s}\in\mathcal{L}_{N}$. Hence, by applying the inequality from (15), we obtain
$\mathcal{B}(\tilde{s})V^{*}$ $<\mathcal{B}(\tilde{s})\mathcal{C}(a)+\alpha\mathcal{B}(\tilde{s})\mathcal{T}(a)V^{*}\quad\forall a\in A,$
$\implies\mathcal{B}(\tilde{s})V^{*}$ $<V_{AS;0}(\tilde{s}).$ (16)
Thus, for all $\tilde{s}\in\mathcal{L}_{N}$, where $N\geq|\mathcal{S}|-2$, applying (16) to the definition of $V_{AS;0}(\tilde{s})$ in (14), we obtain
$V_{AS;0}(\tilde{s})<\mathcal{B}(\tilde{s})\mathcal{C}(a)+\alpha V_{AS;0}(\tilde{s}_{a})\quad\forall a\in A$ (17)
Now let $i^{\prime}$ be the state reached after taking a blind step with action $a$ from state $i\in\mathcal{L}_{N}$. Then, it immediately follows from (17) that
$Z(i)+\alpha^{N}V_{AS;0}(i)<Z(i)+\alpha^{N}\left(\mathcal{B}(i)\mathcal{C}(a)+\alpha V_{AS;0}(i^{\prime})\right)\\ =Z(i^{\prime})+\alpha^{N+1}V_{AS;0}(i^{\prime}).$
Thus
$\min_{i\in\mathcal{L}^{j}_{N}}\left(Z(i)+\alpha^{N}V_{AS;0}(i)\right)$ $<\min_{i\in\mathcal{L}^{j}_{N+1}}\left(Z(i)+\alpha^{N+1}V_{AS;0}(i)\right)$
$\epsilon_{N+1}$ $<\epsilon_{N}.$
To prove $\epsilon_{N+1}\leq\epsilon_{N},$ for a general MDP, we simply replace the strict inequalities in (16) and (17) with non-strict ones. ∎
Appendix B Inventory Management Case Study
This example is adapted from (Bradley et al., 1977). We consider an inventory with a capacity of $3$ units. The demand for items is either $1$ or $2$ units, each with probability 1/2 at every step (month). The production cost for an item is \$$\$1000$ per unit, while the selling price stands at \$$\$2000$ per unit, ensuring a profit of \$$\$1000$ units per sale. We consider a holding cost of \$$\$500$ on each month for each remaining item in the inventory by month-end. [^3] Holding cost is evaluated based on the no. of remaining items of the inventory at the end of the month after meeting the demand. Furthermore, consider a sensing cost of either \$$\$200$ or \$$\$64$ for observing the remaining items in the inventory (the state), and we aim to maximize the discounted profit with the discounting factor $\alpha=0.8$.
Our main takeaways are as follows. For sensing cost \$200, our results are shown in Figure 5. In this case, we see that the heuristic policy (Section 3) performs quite close to the optimal policy (judging by the bound on sub-optimality gap) and the optimal policy for the truncated MDP $\mathcal{M}_{k,N}$ outperforms it only after $N\geq 7$. However, the conditions of Theorem 4.4 are not satisfied over the depths $N$ we were able to compute for (recall that the computational complexity of solving $\mathcal{M}_{k,N}$ grows exponentially in $N$). This is consistent with the results in Figure 5; we continue to see small cost benefits from increasing the threshold on the number of blind actions allowed.
For the lower sensing cost of \$64, our results are shown in Figure 6. In this case, we see that the heuristic policy (Section 3), which does provide an improvement over always sensing, is in fact optimal for $\mathcal{M}_{k}$. Moreover, the optimal policy for $\mathcal{M}_{k,1}$ is also found to be optimal for $\mathcal{M}_{k}$. However, the condition of Theorem 4.4 is only satisfied at $N=7.$
Figure 5: Inventory management case study with sensing cost \$$\$200$
Figure 6: Inventory management case study with sensing cost \$$\$64$